You are here: Foswiki>Commissioning Web>Logs>23Jan013_Log (24 Jan 2013, JuanGuerra)Edit Attach

UT 20130123 AO#1

Observer:Esposito,S(Arcetri),Bussoni,L(Arcetri),Puglisi,A(Arcetri),Pinna,E(Arcetri),JCGuerra(LBTO),JChristou(LBTO),DMiller(LBTO)
Telescope Operator: Steve..(LBTO)
SW Support: Cox,D (Tucson)
AO Support: JCGuerra(LBTO),JChristou(LBTO),DMiller(LBTO))
Telescope Support: Kevin N(LBTO)
Instrument Support: Tim S (LBTO)

Summary

AGW#2 is moved to the lab to investigate the problem of the intermittent power failure of the Wavefront sensor Unit

Issues: -Intermitent power failure of the W-Unit. -TT control. -A connector in the Postdam cooling plate is almost out. This has to be replaced before the unit is back to the telescope and to allow the wavefront sensor unit to be tested at low temperature.

Detail

The AGW#2 is moved to the lab in the morning (about 10:30am). The intermittent power failure could have three of four candidates and to investigate the failure the right electronic box is dismount. The extraction of the box is a not complicate procedure, but the opening of the box is a procedure that it has to be consider with a high risk due to the way that the electronic is packed and all the exposued connectors. The box are opened and it has been found that one of the power supply (XP) which gives an output of different voltages (12,5 and 3) to several devices which one of those is the TS8. The 12 voltage output was a bit low , it was giving 11.57. When the Unit is working and probably in some temperature conditions the power load increase and with this increase will cause a lost of power of the TS8 and the whole AO system can not be power on.

So, it has been found a good candidate.

The spares power supply and ts8 were put in the freezer to make a funtional cold test with a successa at low temperature. But once the box was open with the access to the power suppy it ws identify three trimmer and we could use the trimmer to increase the voltage in the 12 voltage channel. The trimmer was moved a bit to give a 12.7 output voltage. The electronic is assembly and tested in the bench with a total ok, but to test with all the devices the box has to be pu it back to the W Unit.

The unit is out it back and in the first try we hear a bad sound "Pufff" this means, something is already burn.

The box is extracted again and opened, looking the devices and to understand from where this "puff" come from, it ws found one electronic board in the PI was burnt., The PI is replaced with the spare and tested in the bench . To get a full test of the TipTilt, the box has to put it back to the W Unit.

The test with the devices shows that the TT is not working properly in one of the channels, After making some tests with the unit connected and swapping channels, we have found that probably the errors are coming from; the connector of the channel 2 or it could be the strain gauge board.

The connector seems ok.

The box is removed and the strain gauge TT board is replaced with the spare. To test the TT the box is mounted back to the W Unit. After the control test of all the channel the problem of the channel 2 is gone ... !!!! Happy... !!!.

With the Unit repaired, a test of all devices is done without the ccd47. To have a full operation system test the AGW has to bee cold down. To test the unit at low temperature the broken connect in the Postdam AGW part has to be replaced.

-- JuanGuerra - 24 Jan 2013
Topic revision: r2 - 24 Jan 2013, JuanGuerra
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback